OpenStudy (anonymous):

An airline ramp with a base of 50 feet makes an angle of 27° with the ground. I need to find out how long will a person walk on the ramp.

OpenStudy (mathstudent55):

I think this is what the question means. We are looking for x. 

OpenStudy (mathstudent55):

We need the trig function that relates the adjacent leg to the hypotenuse. 

OpenStudy (mathstudent55):

Recalling: SOHCAHTOA, we can use the cosine: \(\cos \theta = \dfrac{adj}{hyp} \) \(\cos 27^o = \dfrac{50 ~ft}{x} \) \(x = \dfrac{50~ft}{\cos 27^o} \) Just do the division to find x, the length of the ramp.